OpenStudy (anonymous):

please i need help i will fan and give medal and write good 'bout you on you're prof.

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

hi

OpenStudy (pixiedust1):

WELCOME TO OPEN STUDY!

OpenStudy (anonymous):

Becca's family has 4 gallons of milk. They drink gallon per day. How long will their milk last? A. days B. 4 days C. days D. days sorry they drink 5/8 gollon per day

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

@PixieDust1 plz help me im asking

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

what are some of the answers

OpenStudy (anonymous):

Becca's family has 4 gallons of milk. They drink gallon per day. How long will their milk last? A. days B. 4 days C. days D. days sorry they drink 5/8 gallons per day

OpenStudy (anonymous):

Bah, the site glitched out on me. Anyhow, you know the number of gallons and how many gallons the family drinks per day. Dividing the gallons by the gallons per day gives you the number of days.\[4 gallons \div \frac{ 5/8 gallon }{ day } = 4 \div \frac{ 5 }{ 8 } = 4*\frac{ 8 }{ 5 } = 6.4 days\]This makes sense, since the less milk they drink per day, the more days the milk will last.

OpenStudy (anonymous):

sorry let me put it in Becca's family has 4 gallons of milk. They drink 5/8 gallon per day. How long will their milk last? A. 6 2/5 B. 4 days C. 2 1/2 D. 4 5/8

OpenStudy (anonymous):

@Krissy1997

OpenStudy (anonymous):

a

OpenStudy (anonymous):

a is the answer

OpenStudy (anonymous):

dude.... your not to tell the answer.... your supose to help them find it them self

OpenStudy (anonymous):

and its not A

OpenStudy (anonymous):

kay thnx @longboardman15

OpenStudy (anonymous):

5/8=0.625 0.625*4=2.5

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Er, no. The family drinks 0.625 gallons of milk per day. 4 gallons will last 6.4 days, not 2.5 days. I'll type up an explanation and post it in a few minutes.

OpenStudy (anonymous):

then what is the answer i just need the answer please

OpenStudy (anonymous):

i just told you

OpenStudy (anonymous):

first you do 5/8

OpenStudy (anonymous):

that equals 0.625

OpenStudy (anonymous):

then you do 0.625*4

OpenStudy (anonymous):

and then you get 2.5

OpenStudy (anonymous):

Krissy, you need to divide by 0.625, not multiply.

OpenStudy (anonymous):

i think its 6 2/5 right ?

OpenStudy (anonymous):

?????????????

OpenStudy (anonymous):

ahhhhhhhhhhhhh to mush mathhhhhhhhhhhhhhhh

OpenStudy (anonymous):

exactly i hate math

OpenStudy (anonymous):

i good at it but dang

OpenStudy (anonymous):

A rate is measured in how much of something happens per some unit of time. For example, speed limits are measured in miles per hour. 30 MPH means that for every hour you travel, you go thirty miles. If you drive for three hours, you'll go 90 miles. I used a similar method to solve your problem. The family drinks 5/8 gallons of milk per day (0.625 gallons/day). After two days, they will have drunk 10/8 gallons of milk (1.25 gallons). After four days, they'll have gone through 20/8 gallons of milk (2.50 gallons). To get how long four gallons of milk will last, you need to divide the amount they have by how much they drink per day. 4 / 0.625 = 6.4 days This makes sense, since the less they drink per day, the longer it will last. For comparison: One gallon per day: 4 / 1 = 4 days Two gallons per day: 4 / 2 = 2 days One half gallon per day: 4 / 0.5 = 8 days One quarter gallon per day: 4 / 0.25 = 16 days

OpenStudy (anonymous):

By the way, 6.4 days = 6 2/5 days. The former is in decimal form; the latter is in the improper fraction form.

OpenStudy (anonymous):

just someone please just tell me the answer

OpenStudy (anonymous):

The answer's been laid out thrice in my responses. Give 'em a read, perhaps?

OpenStudy (anonymous):

please just tell me the answer

OpenStudy (anonymous):

"By the way, 6.4 days = 6 2/5 days" That would be answer choice A. It's highly important that you understand how this works, however, since you can't ask for help on quizzes and tests. I can try a few other ways to explain it if need be.

OpenStudy (anonymous):

yes i already said that the answer is a

OpenStudy (anonymous):

@HELPMEplznthnx the answer is a